OpenStudy (anonymous):

What is the overall cell potential for this redox reaction? Ba + Ca2+ → Ba2+ + Ca Given: Ba2+ + 2e− → Ba(s) –2.90 and Ca2+ + 2e− → Ca(s) –2.87

OpenStudy (rushwr):

Cell potential = potential of the cathode - potential of the anode First you have to figure out what is the anode and cathode. Anode has a higher reducing potential because anode is where oxidation occurs -2.90 is lesser than -2.87 Therefore the anode should be the one with the -2.90 potential and the cathode is the one with -2.87 Cell potential = -2.87 - (-2.90 ) = -2.87 + 2.90 = 0.03 V

OpenStudy (cuanchi):

Ba + Ca2+ → Ba2+ + Ca Ba → Ba2+ + 2e- oxid +2.90 V (reverse the reaction change the sign) Ca2+ + 2e- → Ca red -2.87 V -------------------------------------- Ba + Ca2+ → Ba2+ + Ca E = +2.90V - 2.87 V = 0.03 V